OpenStudy (bossimbacon):

Shawn ran 5 miles more than Alex ran. The sum of their distances is 19 miles. How far did Alex run? The domain of the solution is {7, 12, 14, 24}. Grading Rubric 2 points: Choose a letter to represent your unknown value 2 points: Write an algebraic expression for each person using only one variable 2 points: Write an equation using the two algebraic expressions 2 points: Plug in each value from the domain into the equation 2 points: Find the distance Alex ran Can anyone help me pls?

OpenStudy (haseeb96):

what do u think so, what is the correct answer ?

OpenStudy (bossimbacon):

well I dont quite understand the question and theres many questions to it could you guide me through it? @Haseeb96

jimthompson5910 (jim_thompson5910):

`Choose a letter to represent your unknown value` which letter do you want to pick?

OpenStudy (bossimbacon):

any letter will be fine?

jimthompson5910 (jim_thompson5910):

yeah, usually x works

OpenStudy (bossimbacon):

ok then x

jimthompson5910 (jim_thompson5910):

`Shawn ran 5 miles more than Alex ran` if Alan ran x miles (some unknown number of miles), then how many did Shawn run? in terms of x

OpenStudy (bossimbacon):

s5+a?

OpenStudy (bossimbacon):

i dont know haha

jimthompson5910 (jim_thompson5910):

where are you getting s and a?

jimthompson5910 (jim_thompson5910):

let's use actual numbers say Alan ran 7 miles how many miles did Shawn run?

OpenStudy (bossimbacon):

12

jimthompson5910 (jim_thompson5910):

how did you get that

OpenStudy (bossimbacon):

cause it saids shawn ran 5 more miles than alan

jimthompson5910 (jim_thompson5910):

so you basically add 5 to whatever Alan has if Alan has x, then shawn has x+5

OpenStudy (bossimbacon):

ooooohhh

jimthompson5910 (jim_thompson5910):

adding x to x+5 gives what?

OpenStudy (bossimbacon):

5x?

jimthompson5910 (jim_thompson5910):

1apple + 1apple = 2apples so x+x = 2x

jimthompson5910 (jim_thompson5910):

x + x+5 would turn into 2x+5, right?

OpenStudy (bossimbacon):

yeah

jimthompson5910 (jim_thompson5910):

`The sum of their distances is 19 miles` we just found the total distance to be 2x+5 equate the two 2x+5 = 19 then solve for x

OpenStudy (bossimbacon):

x would be 7 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

and x+5 would be x+5=7+5 = 12 so Alan ran 7 miles, Shawn ran 12 miles 7+12 = 19 checks out

OpenStudy (bossimbacon):

but does this answer each question? 2 points: Choose a letter to represent your unknown value 2 points: Write an algebraic expression for each person using only one variable 2 points: Write an equation using the two algebraic expressions 2 points: Plug in each value from the domain into the equation 2 points: Find the distance Alex ra

jimthompson5910 (jim_thompson5910):

how did you find that x = 7

jimthompson5910 (jim_thompson5910):

did you plug in various x values? or did you isolate x?

OpenStudy (bossimbacon):

plug in

jimthompson5910 (jim_thompson5910):

so you did the `Plug in each value from the domain into the equation` step. Be sure to list every value from that set

OpenStudy (bossimbacon):

oh no i didnt ill make sure to put that in thank you! but what about the other questions?

OpenStudy (bossimbacon):

2 points: Write an algebraic expression for each person using only one variable?

jimthompson5910 (jim_thompson5910):

that's when I said that alan ran x miles shawn ran x+5 miles

OpenStudy (bossimbacon):

Write an equation using the two algebraic expressions

OpenStudy (bossimbacon):

what about this question? Write an equation using the two algebraic expressions

jimthompson5910 (jim_thompson5910):

I wrote the equation above too

OpenStudy (bossimbacon):

@jim_thompson5910 thank you so much! really means alot! I appreciate your patience :)

jimthompson5910 (jim_thompson5910):

you're welcome